# Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21

### v4

@@ Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
I Appendix
Comparison of the theory to simulations
Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius excited by an oscillating dipole located at with a dipole moment of . Top: over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. The circles are the peak widths obtained analytically from compared to the ones calculated from the simulations, with very good agreement. (a),(b),(c),(d),(e) for the resonance conditions: calculated at
respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement. Bottom: over the particle envelope calculated in COMSOL for a dipole located at (same dipole distance as the previous case) with a dipole moment of As predicted the modes disappeared from the spectrum.Figure 5: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 sphere of a radius excited by an oscillating dipole located at with a dipole moment of . Top: calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. (a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement. Figure 6: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 ellipsoid with excited by an oscillating dipole located at with a dipole moment of . Top: Surface integral over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. (a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig. 5, defined in the ellipsoid volume).
We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs. 2,3 [55]. The sphere is of radius and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of since we assume [67, 73, 2], and later on prove, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in . This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, see SM for details. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflects the presence of higher spatial frequencies near the dipole, which excite higher-order modes. We then calculate the peak widths analytically and from the simulations. To demonstrate the mode coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis. In this case we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required.
In Fig. 4 we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows in Fig. 4) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t accounted for it). Clearly, there is remarkable agreement between them. This showcases our important ability to analytically predict the Q factors given by
We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes. While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm). As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles.
We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref. [55]. The sphere is excited by a dipole located at with the dipole moment of
We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. Here, again, the dipole moment is directed along to preserve the angular dependency of the mode in the mode-excitation coefficient. In Fig. 5 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of . Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig. 5. We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement. We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with composed of MoO3 with the permittivity model in Ref. [55]. The ellipsoid is excited by a dipole located at with the dipole moment of
We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. In Fig. 6 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were easily calculated by finding the maxima of . We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement. We note that the isotropic ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following:
As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonance are split, as can be seen in Fig. 6. This differs from the case of spheres where the splitting occurs for modes of the same order. Also, in this case the anisotropy of the material could in principle lead to resonance merging.
Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellispoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
II Supplementary Material
II.1 Correspondence between the derived modes and the radiation patterns
@@
Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres. We utilize the TM vector spherical harmonics (VSH). This field pattern is proportional to the sphere mode outside the inclusion [74], which gives the quasistatic results in the limit of Even though this mode is usually considered for isotropic spheres, the results can be applied to the modes of anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs. We write a TM VSH:
where is the spherical Hankel function of the first kind.
Using the vector identity for
we get for the first term:
where
For the second term we use the identity
to get
Then, we utilize the spherical-harmonic identity
@@
We see that when the second term, which is longitudinal, dominates with scaling, and for
the first term, which is transverse, dominates (actually its first term that scales as
Thus, one can conclude from the superposition of that participate in the
quasistatic anisotropic modes, that the radiation pattern will be of the form:
Please note that which gives
overall scaling of the radiated field and we get the following radiation pattern:
II.2 Biaxial sphere modes as superpositions of spherical harmonics
@@
II.3 Completeness of the solution
We now prove that the derived uniaxial-sphere modes are all the eigenmodes.
Let us assume that there is an additional eigenstate besides the four modes for each .
Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
@@ II.3 Completeness of the solution
If we get which is a uniaxial medium, and we therefore exclude this case.
All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints. As a result, their potential contribution is typically weak. Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
II.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
II.4.1 Theory
We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the Appendix.
Dipole mode
The resonance condition is and we therefore find the maximum of
which occurs for hBN [55] at The mode field inside the particle is given by
2nd order modes
For the first 2nd-order mode we get from the theory
which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
second mode
The derived resonance condition reads
For hBN this is approximately satisfied for
Now we calculate the field by adding the degenerate modes:
3rd order modes
First mode
The resonance condition reads:
For hBN this occurs at
. Adding the degenerate modes we get
To simplify things, we look at the plane
Second mode
The resonance condition is satisfied
for hBN for Adding the modes we obtain:
We plot these field magnitude distributions in Fig. 4.
II.5 Eigenmode expansion for an anisotropic inclusion
Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [74, 67].
Here we expand the quasielectrostatic potential for an anisotropic inclusion.
For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
@@ II.5 Eigenmode expansion for an anisotropic inclusion
where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity.
Now we move one term to the lhs in order to be able to define an
eigenvalue equation.
where . The Green function in this case will be for a medium with an inclusion
with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
@@
We express the electric potential and expand it using the anisotropic
eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
@@
where is the electric potential for the dipole in free space.
All in all we get:
Note that in Fig. 4 there is a very small peak at This peak may be due to an off-resonance uniaxial mode as could be understood from the above expansion when applied to biaxial inclusions. Such an expansion is composed of three sums where the second one corresponds to the off resonance uniaxial-mode contributions. This qualitatively agrees with the potential calculated from the field in the simulation results of the form similarly to the mode constructed by adding the modes. Interestingly, this very small peak is not present in the spectrum of the biaxial ellipsoid, in agreement with our prediction that ellipsoids do not favor uniaxial modes.

### v5

@@ Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments. We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
I Methods
I.1 Correspondence between the derived modes and the radiation patterns
Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres. We utilize the TM vector spherical harmonics (VSH) modes for a general spherical particle given by where
@@ I.1 Correspondence between the derived modes and the radiation patterns
and are the spherical bessel function and spherical Hankel function of the first kind and [79, 20]. Calculating the eigenpermittivity from these modes in the limit of gives the quasistatic eigenvalues. For example, the full electrodynamic mode gives the following eigenpermittivities for a sphere with nm and nm : which are very close to the quasistatic results of Even though these modes are usually considered for isotropic spheres, the results can be applied to anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs. We write a TM VSH outside the sphere:
Using the vector identity for we get for the first term:
For the second term we employ the same identity to get
Then, we utilize the spherical-harmonic identity [80]
@@
We see that when since the overall scaling is of in agreement with the quasistatic analysis. For since and the first term, which is transverse and has scaling as expected, dominates. Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation pattern will be of the form:
where the sum is over the mode orders which participate in the anisotropic-sphere modes. Please note that which gives overall scaling of the radiated field and we get the following radiation pattern:
One can readily see that the mode of a subwavelength sphere gives the dipole radiation pattern of . In addition, the radiation patterns of high-order anisotropic particle modes differ from the isotropic case. Note that while the uniaxial-sphere modes are identical to the isotropic sphere modes, in nearly all practical scenarios, their radiation patterns differ due to the fact that an excitation source will excite distinct superpositions of modes in the two cases.
I.2 Biaxial sphere modes as superpositions of spherical harmonics
@@
I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
I.3.1 Theory
We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the main text.
Dipole mode
The resonance condition is and we therefore find the maximum of which occurs for hBN [54] at The mode field inside the particle is given by
2nd order modes
For the first 2nd-order mode we get from the theory which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
second mode
The derived resonance condition reads
For hBN this is approximately satisfied for Now we calculate the field by adding the degenerate modes:
3rd order modes
First mode
The resonance condition reads:
For hBN this occurs at . Adding the degenerate modes we get
To simplify things, we look at the plane
Second mode
The resonance condition is satisfied for hBN for Adding the modes we obtain:
We plot these field magnitude distributions in Fig. 4.
Supplementary Material
Completeness of the solution
We now prove that the derived uniaxial-sphere modes are all the eigenmodes. Let us assume that there is an additional eigenstate besides the four modes for each . Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
@@ I.3.1 Theory
If we get which is a uniaxial medium, and we therefore exclude this case. All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints. As a result, their potential contribution is typically weak. Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
Eigenmode expansion for an anisotropic inclusion
Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [79, 67]. Here we expand the quasielectrostatic potential for an anisotropic inclusion. For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
@@
where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity. Now we move one term to the lhs in order to be able to define an eigenvalue equation.
where . The Green function in this case will be for a medium with an inclusion with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
@@
We express the electric potential and expand it using the anisotropic eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
@@
where is the electric potential for the dipole in free space. All in all we get: